OpenStudy (anonymous):

need help, im really confused which of the following equations is the transaltion 2 units down of the graph of y=(x)?

OpenStudy (whpalmer4):

Translation up and down the y-axis is accomplished by adding or subtracting to the result of the function. For example, \(y=x^2+2\) is the graph of \(y=x^2\) shifted up by 2 units. \(y=x^2-2\) is the graph of \(y=x^2\) shifted down by 2 units.